# Neuro-Symbolic Entity Alignment via Variational Inference

### v1

@@ 7 Conclusions
In this work, we introduced NeuSymEA, a neuro-symbolic framework for entity alignment that integrates symbolic and neural methods within a variational EM framework. NeuSymEA addresses challenges such as substructure heterogeneity, sparsity, and uncertainty. Empirical results show it outperforms existing approaches while offering interpretable alignment predictions via a path-ranking-based explainer. This demonstrates NeuSymEA’s potential to advance knowledge fusion by enabling effective and uncertainty-aware entity alignment.
@@ A.1 Notations
${\mathcal{E}},{\mathcal{E}}^{\prime}$ The sets of entities in ${\mathcal{G}}$ and ${\mathcal{G}}^{\prime}$, respectively
${\mathcal{R}},{\mathcal{R}}^{\prime}$ The sets of relations in ${\mathcal{G}}$ and ${\mathcal{G}}^{\prime}$, respectively
@@
${\mathcal{H}}$ Set of unobserved entity pairs, i.e., ×\${\mathcal{E}}\times{\mathcal{E}}^{\prime}\backslash{\mathcal{O}}$
@@
Table 3: Notations
A.2 Pseudo-code of Explainer
@@ 1. Search Reachable Anchor Pairs within Max Depth $\mathcal{L}$
  $RN\leftarrow\text{BFS}(e_{q},\mathcal{G},\mathcal{L})$ /* Search reachable neighbors of $e_{q}$ using breadth-first search, max depth $\mathcal{L}$/
  $RN^{\prime}\leftarrow\text{BFS}(e_{q}^{\prime},\mathcal{G}^{\prime},\mathcal{L})$ /* Search reachable neighbors of $e_{q}^{\prime}$ using breadth-first search, max depth $\mathcal{L}$/
@@ B.1 Dataset statistics
The DBP15K dataset is designed for cross-lingual knowledge graph alignment and comes in two versions: the full version and the condensed version. The full version of DBP15K includes comprehensive data across three language pairs. The condensed version eliminates the less frequent entities and their connected triples. Statistics of both versions are presented below.
@@
Table 4: Data statistics of the full DBP15K dataset.
@@
Table 5: Data statistics of the condensed DBP15K dataset.
B.2 Comprehensive Results with different ratio of training data
Below we present the comprehensive performance evaluation with various ratios of pairs as training data. In these 3x3 subfigures, each row shows the results of the same dataset, and each column shows the results of the same metric. To summarize, we have the following observations from Figure 5:
Figure 5: Alignment performance with varying percentages of pairs as training data.
  • Neural-symbolic models, particularly our proposed NeuSymEA (NeuSymEA-L and NeuSymEA-D), consistently outperform other approaches across all datasets and metrics. These models demonstrate remarkable robustness, maintaining high performance even with limited training data (as low as 1%). Their success highlights the effectiveness of combining neural and symbolic approaches in entity alignment tasks, offering a significant advantage over traditional neural or purely symbolic methods.
  • The performance of all models generally improves as the percentage of training data increases, but the rate of improvement varies significantly. Traditional neural models like GCNAlign and AlignE show the steepest improvement curves, indicating their heavy reliance on large training datasets. In contrast, neural-symbolic models, especially our proposed NeuSymEA, demonstrate high performance even with minimal training data, showcasing their efficiency and potential for low-resource scenarios.
  • The experiments reveal varying levels of difficulty in entity alignment across different language pairs. Japanese-English alignment generally shows the highest performance across models, followed by French-English, while Chinese-English proves to be the most challenging. This variance underscores the importance of considering language-specific characteristics in entity alignment tasks and suggests that some language pairs may require more sophisticated approaches or additional resources for effective alignment. The observed differences might be attributed to factors such as linguistic distance, writing system differences, or the availability and quality of pre-existing resources for each language pair. These findings emphasize the importance of developing flexible models that can adapt to the specific challenges posed by different language combinations.
  • The exceptional performance of NeuSymEA across various scenarios – different language pairs, metrics, and data availability levels – points to their robustness and scalability. This consistency suggests that these models might be more easily adaptable to new languages or domains without requiring extensive retraining or modification. Such robustness is crucial for developing general-purpose entity alignment systems that can be deployed across diverse linguistic and domain-specific contexts. Additionally, its ability to perform well with limited data indicates potential for scalability to a wider range of languages and domains, including those with limited resources, which could significantly expand the reach and applicability of cross-lingual knowledge integration technologies.
B.3 Parameter analysis
We present the hit@1 performance of NeuSymEA across three datasets, varying hyperparameters, illustrated by a three-dimensional graph. The threshold hyperparameter $\delta$ is explored within the set {0.6, 0.7, 0.8, 0.9, 0.95, 0.98, 0.99}, while the number of EM iterations ranges from 1 to 9. Performance levels are indicated using a colormap. Performance sensitivity analysis in Figure 6 reveals that for all datasets, performance generally improves as the iteration increases. On the other hand, the performance is less sensitive to the threshold $\delta$.

### v2

@@ 8 Conclusions
We presented NeuSymEA, a unified and extensible neuro-symbolic framework for entity alignment. By unifying neural and symbolic reasoning, NeuSymEA addresses the challenges of substructure heterogeneity, sparsity, and uncertainty in real-world KGs. Empirical results demonstrate NeuSymEA’s clear improvements over baselines and robustness under limited resources. By delivering interpretable alignment predictions with uncertainty scores, NeuSymEA advances knowledge fusion, enabling effective and trustworthy entity alignment.
@@ A.1 Notations
$\mathcal{E}$$\mathcal{E}$$^{\prime}$ The sets of entities in $\mathcal{G}$} and $\mathcal{G}$$^{\prime}$, respectively
$\mathcal{R}$$\mathcal{R}$$^{\prime}$ The sets of relations in $\mathcal{G}$} and $\mathcal{G}$$^{\prime}$, respectively
@@
$\mathcal{H}$} Set of unobserved entity pairs, i.e., ${\mathcal{E}}\times{\mathcal{E}}^{\prime}\backslash{\mathcal{O}}$
@@
A.2 Complexity analysis of the symbolic reasoning
In the following, we present the analysis of runtime complexity and parameter complexity one by one.
A.2.1 Runtime complexity
In variational inference, the process of learning and inferring long rules (Equation 7) is simplified by decomposing them into unit-length rules (Equation 8). Consequently, rule weight learning (Equation 10) is only conducted for unit-length rules. The inference process for an $L$-length rule is then estimated by iteratively applying inference steps with unit-length rules (Equation 9) for $L$ iterations. This strategy effectively avoids the exponential search space associated with longer rules, making the computational complexity of the inference linear with respect to the rule length $L$.
Each iteration of reasoning with unit-length rules comprises an inference step (Equation 9) and a rule-weight learning step (Equation 10). These steps require computing the matching probability for all possible entity pairs and relation pairs, respectively. As a result, the computational complexity of the inference step and the weight updating step are $O(|\mathcal{E}||\mathcal{E}^{\prime}|)$ and $O(|\mathcal{R}||\mathcal{R}^{\prime}|)$, respectively.
Thus, the total computational complexity for reasoning with an $L$-length rule is $O(L\cdot(|\mathcal{E}||\mathcal{E}^{\prime}|+|\mathcal{R}||\mathcal{R}^{\prime}|))$. Given that entity sizes are typically much larger than relation sizes, this complexity can be approximated as $O(L\cdot|\mathcal{E}||\mathcal{E}^{\prime}|)$.
Notably, the computations involved in Equation 9 and Equation 10 can be accelerated through parallel processing, which we have implemented. This optimization reduces the runtime complexity to $O\left(L\cdot\frac{|\mathcal{E}||\mathcal{E}^{\prime}|}{n}\right)$, where $n$ represents the number of CPU cores available for parallelization.
A.2.2 Parameter complexity
The total number of alignment probabilities for all entity pairs is $|{\mathcal{E}}||{\mathcal{E}}^{\prime}|$, which is large when the entity sizes increase. We adopt a lazy inference strategy to enhance parameter efficiency. This strategy involves only saving the alignment probabilities of the most probable alignments:
$\left\{p_{w}(v_{(e_{i},e_{i}^{\prime})}),|,e_{i}\in\mathcal{E},e_{i}^{\prime}\in\mathcal{E}^{\prime},p_{w}(v_{(e_{i},e_{i}^{\prime})})=\max\left(\max_{e\in\mathcal{E}}p_{w}(v_{(e,e_{i}^{\prime})}),\max_{e^{\prime}\in\mathcal{E}^{\prime}}p_{w}(v_{(e_{i},e^{\prime})})\right)\right\}$ (12)
Probabilities of other entity pairs can be inferred from these saved alignment probabilities using Equation 9. In this way, parameter complexity is reduced to $O(\max\left(|\mathcal{E}|+|{\mathcal{E}}^{\prime}|)\right)$.
A.3 Pseudo-code of Explainer
@@ 1. Search Reachable Anchor Pairs within Max Depth $\mathcal{L}$
  $RN\leftarrow\text{BFS}(e_{q},\mathcal{G},\mathcal{L})$ /* Search reachable neighbors of $e_{q}$ using breadth-first search, max depth $\mathcal{L}$ */
  $RN^{\prime}\leftarrow\text{BFS}(e_{q}^{\prime},\mathcal{G}^{\prime},\mathcal{L})$ /* Search reachable neighbors of $e_{q}^{\prime}$ using breadth-first search, max depth $\mathcal{L}$ */
@@ B.1 Dataset statistics
The DBP15K dataset, designed for cross-lingual knowledge graph alignment, has two versions: full and condensed. The original full version resembles real-world knowledge graphs, including comprehensive data across three language pairs with many sparsely connected low-degree entities. The condensed version, derived by JAPE and adopted by later methods (GCNAlign, RREA, Dual-AMN, LightEA), removes these low-degree entities and their connected triples to create a smaller (and higher average degree) dataset suitable for GCN-based methods. Detailed information about the two dataset versions can be found in the "dataset" section of readme in JAPE’s official implementation [^1] https://github.com/nju-websoft/JAPE.
The dataset statistics of them are shown in Table˜5 and Table˜6. In this paper, we adopt both versions of DBP15K for comprehensive evaluation. Specifically, the full dataset is sparser and larger in scale in scale due to the inclusion of low-degree entities, thus suitable for evaluating the models’ robustness to sparsity and large scale.
Table 5: Data statistics of the full DBP15K dataset.
@@
Table 6: Data statistics of the condensed DBP15K dataset.
@@
B.2 Efficiency analysis
To provide a comprehensive understanding of the computational efficiency of our model, we report the runtime and memory usage during the experiments. The results, as summarized in Table Table˜7, demonstrate that our model achieves efficient performance with a runtime of 15 minutes, a memory consumption of 868 MB, and a GPU memory usage of 4.33 GB. These metrics highlight the practicality of our approach in terms of resource utilization. Hardware configurations of the experiments are presented in Table˜8.
Table 7: Runtime and Memory Usage
Runtime Memory GPU Memory
15 minutes 868 MB 4.33 GB
Table 8: Machine configuration.
Component Specification
GPU NVIDIA GeForce RTX 3090
CPU Intel(R) Xeon(R) Silver 4214R CPU @ 2.40GHz